# Disjunctive Benders Decomposition

### v1

@@ 1.5 Notation and Organization of the Paper
This paper is organized as follows: Section 2 presents the proposed approach, starting with the formal definition of the disjunction in the Benders reformulation in Section 2.1, followed by the development of a cut-generating problem in Section 2.2, an oracle for solving this problem in Section 2.3, and the disjunctive Benders decomposition incorporating the proposed oracle in Section 2.4. Section 3 discusses further advancements for mixed-binary linear programs. Section 4 compares the proposed approach with relevant literature. Section 5 evaluates the effectiveness of the disjunctive Benders decomposition through experimental studies. Finally, Section 6 concludes the paper.
@@ 2.1 Disjunction of the Benders reformulation
Note that $\operatorname{conv}\left((\mathcal{X}\times\mathbb{R})\cap\mathcal{L}\right)\subseteq\operatorname{conv}(\mathcal{P}^{(\phi,\phi_{0})})\subseteq\mathcal{P}\subseteq\mathcal{L}$, as the split conveys partial integrality information. It is known that the second inclusion is strict when at least one vertex $x^{\prime}$ of $\mathcal{P}$ satisfies $\phi_{0}<\phi^{\top}x^{\prime}<\phi_{0}+1$ (see, e.g., [12, Proposition 5.2]). Moreover, when all integer variables are binaries (i.e., $[l_{j},u_{j}]=[0,1]$ for all $j\in\mathcal{I}=\{1,\cdots,\ell\}$), the following holds from the sequential convexification theorem [2]:
@@
2.2 Cut-generating linear program (CGLP)
@@ 2.2.2 Normalization constraints
$\max\{\gamma_{0}-\gamma_{x}^{\top}\hat{x}-\gamma_{t}\hat{t}:\gamma\in\Gamma\}.$
Since the feasible region $\Gamma$ is conic (i.e., $\alpha\gamma\in\Gamma$ for any $\alpha\in\mathbb{R}_{\geq 0}$ and $\gamma\in\Gamma$), Problem (2.2.2) becomes unbounded whenever $(\hat{x},\hat{t})\notin\operatorname{conv}(\mathcal{P}^{(\phi,\phi_{0})})$. To address this issue, it is necessary to impose a normalization constraint on (2.2.2) to ensure boundedness and generate meaningful inequalities.
@@ 2.3 The oracle
Let $(\hat{\omega}_{x},\hat{\omega}_{t},\hat{\omega}_{0})\in\mathbb{R}^{n+1}\times\mathbb{R}_{>0}$. Suppose there exists a hyperplane of the form $\alpha_{x}^{\top}x+\alpha_{t}t\geq\alpha_{0}$ that separates $(x^{\prime},t^{\prime}):=(\hat{\omega}_{x}/\hat{\omega}_{0},\hat{\omega}_{t}/\hat{\omega}_{0})$ from $\mathcal{L}$. Then, the hyperplane $\alpha_{x}^{\top}\omega_{x}+\alpha_{t}\omega_{t}\geq\alpha_{0}\omega_{0}$ separates $(\hat{\omega}_{x},\hat{\omega}_{t},\hat{\omega}_{0})$ from $\mathcal{L}^{\#}$. Conversely, if $(x^{\prime},t^{\prime})\in\mathcal{L}$, meaning no such separating hyperplane exists, then $(\hat{\omega}_{x},\hat{\omega}_{t},\hat{\omega}_{0})\in\mathcal{L}^{\#}$.
@@ 3.2 Cut lifting
The Oracle, i.e., Routine 1 or 2, can be called at any fractional node of a branch-and-bound tree when solving (1.1) in the BenderBnB framework. Each node $n$ of a branch-and-bound tree is associated with a pair of disjoint index sets $(\mathcal{I}_{n,0},\mathcal{I}_{n,1})$, where $\mathcal{I}_{n,l}$ denotes the set of indices of binary variables which values are fixed at $l$ at the node. A cut that is generated at $n$ can be extended to be valid for (1.1) by adjusting coefficients for the variables $x_{j}$ where $j\in\mathcal{I}_{n}:=\mathcal{I}_{n,0}\cup\mathcal{I}_{n,1}$ and the right-hand side. This process is referred to as lifting the cut. In this section, we propose a simple postprocessing step for lifting the disjunctive Benders cut.
@@ 4 Discussion
where $(\mathcal{X}\times\mathbb{R})\cap\mathcal{L}_{j}=\{(x,t_{j}):t_{j}\geq f_{j}(x),\ x\in\mathcal{X}\}$. Both the cut-and-project approach and the Lagrangian cut aims to separate a candidate solution $(\hat{x},\hat{t}_{j})$ of (4.1) from $\operatorname{conv}((\mathcal{X}\times\mathbb{R})\cap\mathcal{L}_{j}))$ whenever $(\hat{x},\hat{t}_{j})\not\in\operatorname{conv}((\mathcal{X}\times\mathbb{R})\cap\mathcal{L}_{j}))$ for each $j\in[N]$. They generate cuts based on two different descriptions of $\operatorname{conv}((\mathcal{X}\times\mathbb{R})\cap\mathcal{L}_{j}))$. Specifically, the cut-and-project approach uses the following description:
@@
where $\mathcal{Q}_{j}:=\{(x,t_{j},y_{j}):t_{j}\geq d_{j}^{\top}y_{j},A_{j}x+B_{j}y_{j}\geq b_{j},x\in\mathcal{X}\}$, and the last equality is from [12, Exercise 3.34]. On the other hand, the Lagrangian cut leverages the following alternative description of $\operatorname{conv}((\mathcal{X}\times\mathbb{R})\cap\mathcal{L}_{j})$:
@@
where $f_{j}^{\prime}$ is a function that equals $f_{j}(x)$ for $x\in\mathcal{X}$ and $\infty$ otherwise; ${f_{j}^{\prime}}^{**}$ is the biconjugate of $f^{\prime}_{j}$, whose epigraph is the closure of the convex hull of the epigraph $f^{\prime}_{j}$ [8, Proposition 1.6.1]. The last equality follows from the closesdness of $\operatorname{conv}(\operatorname{epi}f^{\prime}_{j})$, which follows from the lower semicontinuity of $f^{\prime}_{j}$ and the compactness of $\mathcal{X}$. Since we have
@@ 5 Numerical experiments
As discussed in Section 4, the key advantage of disjunctive Benders decomposition lies in its ability to tackle large problems by strategically leveraging simple, typical oracles. To demonstrate this, we evaluate the proposed method on large-scale instances of the uncapacitated facility location problem (UFLP) and the stochastic network interdiction problem (SNIP)—each presenting distinct structural features, which we describe in detail later. Formal problem formulations are provided in the electronic companion. We assess computational performance through a comparative analysis with conventional Benders decomposition, augmented with structure-exploiting typicalOracles, as well as an off-the-shelf solver.

### v2

@@ 1.5 Notation and Organization of the Paper
This paper is organized as follows: Section 2 presents the proposed approach, starting with the formal definition of the disjunction in the Benders reformulation in Section 2.1, followed by the development of a cut-generating problem in Section 2.2, an oracle for solving this problem in Section 2.3, and the disjunctive Benders decomposition incorporating the proposed oracle in Section 2.4. Section 3 discusses further advancements for mixed-binary linear programs. Section 4 compares the proposed approach with relevant literature. Section 5 evaluates the effectiveness of the disjunctive Benders decomposition through experimental studies. Finally, Section 6 concludes the paper. Omitted proofs are provided in Appendix A.
@@ 2.1 Disjunction of the Benders reformulation
Note that $\operatorname{conv}\left((\mathcal{X}\times\mathbb{R})\cap\mathcal{L}\right)\subseteq\operatorname{conv}(\mathcal{P}^{(\phi,\phi_{0})})\subseteq\mathcal{P}\subseteq\mathcal{L}$, as the split conveys partial integrality information. It is known that the second inclusion is strict when at least one vertex $x^{\prime}$ of $\mathcal{P}$ satisfies $\phi_{0}<\phi^{\top}x^{\prime}<\phi_{0}+1$ (see, e.g., [13, Proposition 5.2]). Moreover, when all integer variables are binaries (i.e., $[l_{j},u_{j}]=[0,1]$ for all $j\in\mathcal{I}=\{1,\cdots,\ell\}$), the following holds from the sequential convexification theorem [2]:
@@
2.2 Cut-generating linear program
@@ 2.2.2 Normalization constraints
$\max\{\gamma_{0}-\gamma_{x}^{\top}\hat{x}-\gamma_{t}\hat{t}:\gamma\in\Gamma\}.$ (2.3)
Since the feasible region $\Gamma$ is conic (i.e., $\alpha\gamma\in\Gamma$ for any $\alpha\in\mathbb{R}_{\geq 0}$ and $\gamma\in\Gamma$), Problem (2.3) becomes unbounded whenever $(\hat{x},\hat{t})\notin\operatorname{conv}(\mathcal{P}^{(\phi,\phi_{0})})$. To address this issue, it is necessary to impose a normalization constraint on (2.3) to ensure boundedness and generate meaningful inequalities.
@@ 2.3 The Oracle
Let $(\hat{\omega}_{x},\hat{\omega}_{t},\hat{\omega}_{0})\in\mathbb{R}^{n_{x}+1}\times\mathbb{R}_{>0}$. Suppose there exists a hyperplane of the form $\alpha_{x}^{\top}x+\alpha_{t}t\geq\alpha_{0}$ that separates $(x^{\prime},t^{\prime}):=(\hat{\omega}_{x}/\hat{\omega}_{0},\hat{\omega}_{t}/\hat{\omega}_{0})$ from $\mathcal{L}$. Then, the hyperplane $\alpha_{x}^{\top}\omega_{x}+\alpha_{t}\omega_{t}\geq\alpha_{0}\omega_{0}$ separates $(\hat{\omega}_{x},\hat{\omega}_{t},\hat{\omega}_{0})$ from $\mathcal{L}^{\#}$. Conversely, if $(x^{\prime},t^{\prime})\in\mathcal{L}$, meaning no such separating hyperplane exists, then $(\hat{\omega}_{x},\hat{\omega}_{t},\hat{\omega}_{0})\in\mathcal{L}^{\#}$.
@@ 3.2 Cut lifting
The Oracle, i.e., Routine 1 or 2, can be called at any fractional node of a branch-and-bound tree when solving (1.1) in the BendersBnB framework. Each node $n$ of a branch-and-bound tree is associated with a pair of disjoint index sets $(\mathcal{I}_{n,0},\mathcal{I}_{n,1})$, where $\mathcal{I}_{n,l}$ denotes the set of indices of binary variables which values are fixed at $l$ at the node. A cut that is generated at $n$ can be extended to be valid for (1.1) by adjusting coefficients for the variables $x_{j}$ where $j\in\mathcal{I}_{n}:=\mathcal{I}_{n,0}\cup\mathcal{I}_{n,1}$ and the right-hand side. This process is referred to as lifting the cut. In this section, we propose a simple postprocessing step for lifting the disjunctive Benders cut.
@@ 4 Discussion
where $(\mathcal{X}\times\mathbb{R})\cap\mathcal{L}_{j}=\{(x,t_{j}):t_{j}\geq f_{j}(x),\ x\in\mathcal{X}\}$. Both the cut-and-project approach and the Lagrangian cut aim to separate a candidate solution $(\hat{x},\hat{t}_{j})$ of (4.1) from $\operatorname{conv}((\mathcal{X}\times\mathbb{R})\cap\mathcal{L}_{j}))$ whenever $(\hat{x},\hat{t}_{j})\not\in\operatorname{conv}((\mathcal{X}\times\mathbb{R})\cap\mathcal{L}_{j}))$ for each $j\in[N]$. They generate cuts based on two different descriptions of $\operatorname{conv}((\mathcal{X}\times\mathbb{R})\cap\mathcal{L}_{j}))$. Specifically, the cut-and-project approach uses the following description:
@@
where $\mathcal{Q}_{j}:=\{(x,t_{j},y_{j}):t_{j}\geq d_{j}^{\top}y_{j},A_{j}x+B_{j}y_{j}\geq b_{j},x\in\mathcal{X}\}$, and the last equality is from [13, Exercise 3.34]. On the other hand, the Lagrangian cut leverages the following alternative description of $\operatorname{conv}((\mathcal{X}\times\mathbb{R})\cap\mathcal{L}_{j})$:
@@
where $f_{j}^{\prime}$ is a function that equals $f_{j}(x)$ for $x\in\mathcal{X}$ and $\infty$ otherwise; ${f_{j}^{\prime}}^{**}$ is the biconjugate of $f^{\prime}_{j}$, whose epigraph is the closure of the convex hull of the epigraph $f^{\prime}_{j}$ [8, Proposition 1.6.1]. The last equality follows from the closedness of $\operatorname{conv}(\operatorname{epi}f^{\prime}_{j})$, which follows from the lower semicontinuity of $f^{\prime}_{j}$ and the compactness of $\mathcal{X}$. Since we have
@@ 5 Numerical experiments
As discussed in Section 4, the key advantage of disjunctive Benders decomposition lies in its ability to tackle large problems by strategically leveraging simple, typical oracles. To demonstrate this, we evaluate the proposed method on large-scale instances of the uncapacitated facility location problem (UFLP) and the stochastic network interdiction problem (SNIP)—each presenting distinct structural features, which we describe in detail later. Formal problem formulations are provided in Appendix B. We assess computational performance through a comparative analysis with conventional Benders decomposition, augmented with structure-exploiting typicalOracles, as well as an off-the-shelf solver.